father is 35 years of age 3 years ago he was 4 time as old his son age of the son this time?
Answer by stanbon(75887)

father is 35 years of age 3 years ago he was 4 time as old his son age of the son this time?
---
f = 35
(f-3) = 4(s-3)
----
Substitute and solve for "s"
(35-3) = 4s-12
32 = 4s-12
4s = 44
s = 11 yrs old (son's age now)
==================================
Cheers,
Stan H.
==============